OpenStudy (anonymous):

cos^2 2x-cos^2 6x=sin4x sin8x prove !

OpenStudy (anonymous):

\[\cos^2(2x) - \cos^2(6x) = \sin(4x) \sin(8x)\]

OpenStudy (anonymous):

thought i had that one.. rewrote it.. turns out i have no idea, im thinking sum to product identity but youve probly gathered that much on your own sorry

OpenStudy (anonymous):

@nabsz.j its ok

OpenStudy (dls):

You can write this as: (cos2x+cos6x)(cos2x-cos6x) Now,use cosA+cosB and cosA-cosB You can solve now?

OpenStudy (anonymous):

i did that and i used the next formula too .. tell me what to do after that !

OpenStudy (dls):

okay so now..

OpenStudy (dls):

You should get this: (2 sin 4x cos 4x) (2 sin 2x cos 2x)

OpenStudy (dls):

So,finally sin 8x sin 4x is the answer :)

OpenStudy (anonymous):

i got (2 cos 4x cos 2x ) (-2 sin 4x sin 2x )

OpenStudy (anonymous):

how'd yu do it @DLS ?

OpenStudy (anonymous):

which sin formua ?

OpenStudy (dls):

sin2x=2sinxcosx

OpenStudy (anonymous):

can you do it .. please , am gettin confused !

OpenStudy (dls):

okay!

OpenStudy (dls):

cos2(2x)−cos2(6x)⟹(cos(2x)−cos(6x))(cos(2x)+cos(6x)) So: Using: cos(C)+cos(D)=2cos(C+D/2)cos(C−D)/2) cos(C)−cos(D)=2sin(C+D/2)sin(D−C/2) (2sin(4x)sin(2x))(2cos(4x)cos(2x))⟹(2sin(2x)cos(2x))⋅(2sin(4x)cos(4x)) ⟹sin(4x)⋅sin(8x)

OpenStudy (anonymous):

okay understood thanks DLS :)

OpenStudy (dls):

One step answer :P cos^2A−cos^2B=sin(A+B)sin(B−A)

OpenStudy (anonymous):

how come (2sin(2x)cos(2x))⋅(2sin(4x)cos(4x))=sin(4x)⋅sin(8x) @DLS